\begin{thebibliography}{30}
\providecommand{\natexlab}[1]{#1}

\bibitem[{Ahmad et~al.(2020)Ahmad, Hussan, and Butt}]{ahmad2020english}
Ayaz Ahmad, Sana Hussan, and Ayesha Butt. 2020.
\newblock English in the domain of legal services: Problems of access in pakistan.
\newblock \emph{Pakistan Journal of Criminology}, 12(1):109--127.

\bibitem[{{Anthropic}(2024)}]{anthropic-docs-models}
{Anthropic}. 2024.
\newblock \href {https://docs.anthropic.com/en/docs/about-claude/models} {{Anthropic Documentation: Claude Models}}.
\newblock Accessed: 2024-10-10.

\bibitem[{Arif et~al.(2024)Arif, Farid, Athar, and Raza}]{arif-etal-2024-uqa}
Samee Arif, Sualeha Farid, Awais Athar, and Agha~Ali Raza. 2024.
\newblock \href {https://aclanthology.org/2024.lrec-main.1497} {{UQA}: Corpus for {U}rdu question answering}.
\newblock In \emph{Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024)}, pages 17237--17244, Torino, Italia. ELRA and ICCL.

\bibitem[{Bai et~al.(2023)Bai, Bai, Chu, Cui, Dang, Deng, Fan, Ge, Han, Huang, Hui, Ji, Li, Lin, Lin, Liu, Liu, and Lu}]{bai_qwen_2023}
Jinze Bai, Shuai Bai, Yunfei Chu, Zeyu Cui, Kai Dang, Xiaodong Deng, Yang Fan, Wenbin Ge, Yu~Han, Fei Huang, Binyuan Hui, Luo Ji, Mei Li, Junyang Lin, Runji Lin, Dayiheng Liu, Gao Liu, and Chengqiang et~al. Lu. 2023.
\newblock \href {https://doi.org/10.48550/arXiv.2309.16609} {Qwen {Technical} {Report}}.
\newblock \emph{arXiv preprint}.
\newblock ArXiv:2309.16609 [cs].

\bibitem[{Banerjee and Lavie(2005)}]{banerjee-lavie-2005-meteor}
Satanjeev Banerjee and Alon Lavie. 2005.
\newblock \href {https://aclanthology.org/W05-0909} {{METEOR}: An automatic metric for {MT} evaluation with improved correlation with human judgments}.
\newblock In \emph{Proceedings of the {ACL} Workshop on Intrinsic and Extrinsic Evaluation Measures for Machine Translation and/or Summarization}, pages 65--72, Ann Arbor, Michigan. Association for Computational Linguistics.

\bibitem[{Barrault et~al.(2023)Barrault, Chung, Meglioli, Dale, Dong, Duquenne, Elsahar, Gong, Heffernan, Hoffman, Klaiber, Li, Licht, and Maillard}]{communication_seamlessm4t:_2023}
Loïc Barrault, Yu-An Chung, Mariano~Cora Meglioli, David Dale, Ning Dong, Paul-Ambroise Duquenne, Hady Elsahar, Hongyu Gong, Kevin Heffernan, John Hoffman, Christopher Klaiber, Pengwei Li, Daniel Licht, and Jean et~al. Maillard. 2023.
\newblock \href {https://doi.org/10.48550/arXiv.2308.11596} {{SeamlessM4T}: {Massively} {Multilingual} \& {Multimodal} {Machine} {Translation}}.
\newblock \emph{arXiv preprint}.
\newblock ArXiv:2308.11596 [cs].

\bibitem[{{Central Intelligence Agency}(2023)}]{cia_world_factbook}
{Central Intelligence Agency}. 2023.
\newblock \href {https://www.cia.gov/the-world-factbook/countries/world/#people-and-society} {{The World Factbook - World: People and Society}}.
\newblock Accessed: 2024-10-12.

\bibitem[{Chen et~al.(2023)Chen, Yao, Zhao, Zhang, Sun, Liu, and Shen}]{EQUALS}
Andong Chen, Feng Yao, Xinyan Zhao, Yating Zhang, Changlong Sun, Yun Liu, and Weixing Shen. 2023.
\newblock \href {https://doi.org/10.1145/3594536.3595159} {Equals: A real-world dataset for legal question answering via reading chinese laws}.
\newblock In \emph{Proceedings of the Nineteenth International Conference on Artificial Intelligence and Law}, ICAIL '23, page 71–80, New York, NY, USA. Association for Computing Machinery.

\bibitem[{Cui et~al.(2023)Cui, Ning, Li, Chen, Yan, Li, Ling, Tian, and Yuan}]{cui_chatlaw:_2023}
Jiaxi Cui, Munan Ning, Zongjian Li, Bohua Chen, Yang Yan, Hao Li, Bin Ling, Yonghong Tian, and Li~Yuan. 2023.
\newblock \href {https://doi.org/10.48550/arXiv.2306.16092} {Chatlaw: {A} {Multi}-{Agent} {Collaborative} {Legal} {Assistant} with {Knowledge} {Graph} {Enhanced} {Mixture}-of-{Experts} {Large} {Language} {Model}}.
\newblock \emph{arXiv preprint}.
\newblock ArXiv:2306.16092 [cs] version: 1.

\bibitem[{Google(2024{\natexlab{a}})}]{google_gemini}
Google. 2024{\natexlab{a}}.
\newblock \href {https://ai.google.dev/gemini-api/docs/models/gemini} {{Gemini} models}.
\newblock Accessed: 2024-10-13.

\bibitem[{Google(2024{\natexlab{b}})}]{google_vertex_ai_embeddings}
Google. 2024{\natexlab{b}}.
\newblock \href {https://cloud.google.com/vertex-ai/generative-ai/docs/model-reference/text-embeddings-api} {Vertex {AI}: {Text} {Embeddings} {API}}.
\newblock Accessed: 2024-10-13.

\bibitem[{Han et~al.(2012)Han, Kamber, and Pei}]{han2012data}
Jiawei Han, Micheline Kamber, and Jian Pei. 2012.
\newblock \emph{Data Mining: Concepts and Techniques}, 3rd edition.
\newblock The Morgan Kaufmann Series in Data Management Systems. Elsevier.

\bibitem[{Kazi and Khoja(2021)}]{kazi_uquad1.0:_2021}
Samreen Kazi and Shakeel Khoja. 2021.
\newblock \href {https://doi.org/10.48550/arXiv.2111.01543} {{UQuAD1}.0: {Development} of an {Urdu} {Question} {Answering} {Training} {Data} for {Machine} {Reading} {Comprehension}}.
\newblock \emph{arXiv preprint}.
\newblock ArXiv:2111.01543 [cs].

\bibitem[{Khan et~al.(2024)Khan, Sharif, and Shahid}]{khan_bridging_2024}
Rana Umair~Abrar Khan, Hassan~Sattar Sharif, and Ali Shahid. 2024.
\newblock \href {https://doi.org/10.55737/qjss.501649280} {Bridging the {Gap}: {The} {Transformative} {Role} of {Technology} in {Enhancing} {Access} to {Justice} in {Pakistan}}.
\newblock \emph{Qlantic Journal of Social Sciences}, 5(1):259--267.

\bibitem[{Krasadakis et~al.(2024)Krasadakis, Sakkopoulos, and Verykios}]{electronics13030648}
Panteleimon Krasadakis, Evangelos Sakkopoulos, and Vassilios~S. Verykios. 2024.
\newblock \href {https://doi.org/10.3390/electronics13030648} {A survey on challenges and advances in natural language processing with a focus on legal informatics and low-resource languages}.
\newblock \emph{Electronics}, 13(3).

\bibitem[{Lewis et~al.(2020)Lewis, Perez, Piktus, Petroni, Karpukhin, Goyal, K\"{u}ttler, Lewis, Yih, Rockt\"{a}schel, Riedel, and Kiela}]{NEURIPS2020_6b493230}
Patrick Lewis, Ethan Perez, Aleksandra Piktus, Fabio Petroni, Vladimir Karpukhin, Naman Goyal, Heinrich K\"{u}ttler, Mike Lewis, Wen-tau Yih, Tim Rockt\"{a}schel, Sebastian Riedel, and Douwe Kiela. 2020.
\newblock \href {https://proceedings.neurips.cc/paper_files/paper/2020/file/6b493230205f780e1bc26945df7481e5-Paper.pdf} {Retrieval-augmented generation for knowledge-intensive nlp tasks}.
\newblock In \emph{Advances in Neural Information Processing Systems}, volume~33, pages 9459--9474. Curran Associates, Inc.

\bibitem[{Liu et~al.(2024)Liu, Zhu, Ai, Liu, and Wu}]{liu_ledqa_2024}
Bulou Liu, Zhenhao Zhu, Qingyao Ai, Yiqun Liu, and Yueyue Wu. 2024.
\newblock {LeDQA}: {A} {Chinese} {Legal} {Case} {Document}-based {Question} {Answering} {Dataset}.
\newblock In \emph{Proceedings of the CIKM 2024 Conference}.

\bibitem[{Mistral(2024)}]{mistral_embeddings}
Mistral. 2024.
\newblock \href {https://docs.mistral.ai/getting-started/models/models_overview/} {Models overview}.
\newblock Accessed: 2024-10-13.

\bibitem[{{National Assembly of Pakistan}(2019)}]{constitution_of_pakistan_urdu}
{National Assembly of Pakistan}. 2019.
\newblock \href {https://na.gov.pk/uploads/documents/1575018297_955.pdf} {Constitution of pakistan (in urdu)}.
\newblock Online.
\newblock Accessed: 2024-10-12.

\bibitem[{{OpenAI}(2024)}]{openai_models}
{OpenAI}. 2024.
\newblock \href {https://platform.openai.com/docs/models} {{OpenAI Platform Documentation: Models}}.
\newblock Accessed: 2024-10-10.

\bibitem[{{Pakistan Bureau of Statistics}(2023)}]{pbs_population_2023}
{Pakistan Bureau of Statistics}. 2023.
\newblock \href {https://www.pbs.gov.pk/sites/default/files/population/2023/tables/national/table_11.pdf} {{Population by 5-year age group and sex (2023): National Table 11}}.
\newblock Accessed: 2024-10-12.

\bibitem[{{Pakistan National Assembly}(2023)}]{constitution_of_pakistan_english}
{Pakistan National Assembly}. 2023.
\newblock \href {https://na.gov.pk/uploads/documents/63ea176f52421_610.pdf} {The constitution of the islamic republic of pakistan}.
\newblock Accessed: 2024-10-12.

\bibitem[{Post(2018)}]{post2018clarityreportingbleuscores}
Matt Post. 2018.
\newblock \href {https://arxiv.org/abs/1804.08771} {A call for clarity in reporting bleu scores}.
\newblock \emph{Preprint}, arXiv:1804.08771.

\bibitem[{Rahman et~al.(2023)Rahman, Ghosh, and Arora}]{rahman_utrnet_2023}
Abdur Rahman, Arjun Ghosh, and Chetan Arora. 2023.
\newblock \href {https://doi.org/10.1007/978-3-031-41734-4_19} {{UTRNet}: High-resolution urdu text recognition in printed documents}.
\newblock In \emph{Document Analysis and Recognition - ICDAR 2023}, pages 305--324, Cham. Springer Nature Switzerland.

\bibitem[{Rajpurkar et~al.(2018)Rajpurkar, Jia, and Liang}]{rajpurkar_know_2018}
Pranav Rajpurkar, Robin Jia, and Percy Liang. 2018.
\newblock \href {https://doi.org/10.48550/arXiv.1806.03822} {Know {What} {You} {Don}'t {Know}: {Unanswerable} {Questions} for {SQuAD}}.
\newblock \emph{arXiv preprint}.
\newblock ArXiv:1806.03822 [cs].

\bibitem[{Rajpurkar et~al.(2016)Rajpurkar, Zhang, Lopyrev, and Liang}]{rajpurkar-etal-2016-squad}
Pranav Rajpurkar, Jian Zhang, Konstantin Lopyrev, and Percy Liang. 2016.
\newblock \href {https://doi.org/10.18653/v1/D16-1264} {{SQ}u{AD}: 100,000+ questions for machine comprehension of text}.
\newblock In \emph{Proceedings of the 2016 Conference on Empirical Methods in Natural Language Processing}, pages 2383--2392, Austin, Texas. Association for Computational Linguistics.

\bibitem[{Varghese and M.(2024)}]{10533619}
Rejin Varghese and Sambath M. 2024.
\newblock \href {https://doi.org/10.1109/ADICS58448.2024.10533619} {Yolov8: A novel object detection algorithm with enhanced performance and robustness}.
\newblock In \emph{2024 International Conference on Advances in Data Engineering and Intelligent Computing Systems (ADICS)}, pages 1--6.

\bibitem[{Wang(2022)}]{wang2022modernquestionansweringdatasets}
Zhen Wang. 2022.
\newblock \href {https://arxiv.org/abs/2206.15030} {Modern question answering datasets and benchmarks: A survey}.
\newblock \emph{Preprint}, arXiv:2206.15030.

\bibitem[{Xue et~al.(2021)Xue, Constant, Roberts, Kale, Al-Rfou, Siddhant, Barua, and Raffel}]{xue-etal-2021-mt5}
Linting Xue, Noah Constant, Adam Roberts, Mihir Kale, Rami Al-Rfou, Aditya Siddhant, Aditya Barua, and Colin Raffel. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.naacl-main.41} {m{T}5: A massively multilingual pre-trained text-to-text transformer}.
\newblock In \emph{Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies}, pages 483--498, Online. Association for Computational Linguistics.

\bibitem[{Zhong et~al.(2019)Zhong, Xiao, Tu, Zhang, Liu, and Sun}]{zhong_jec-qa:_2019}
Haoxi Zhong, Chaojun Xiao, Cunchao Tu, Tianyang Zhang, Zhiyuan Liu, and Maosong Sun. 2019.
\newblock \href {https://doi.org/10.48550/arXiv.1911.12011} {{JEC}-{QA}: {A} {Legal}-{Domain} {Question} {Answering} {Dataset}}.
\newblock \emph{arXiv preprint}.
\newblock ArXiv:1911.12011 [cs].

\end{thebibliography}

\pdfoutput=1
\documentclass[11pt]{article}
\usepackage[preprint]{acl}
\usepackage{times}
\usepackage{latexsym}
\usepackage[T1]{fontenc}
\usepackage[utf8]{inputenc}
\usepackage{microtype}
\usepackage{inconsolata}
\usepackage{graphicx}
\usepackage{booktabs}
\usepackage{array}
\usepackage{fancyhdr}
\usepackage{hyperref}
\usepackage{tcolorbox}
\usepackage{lipsum}
\usepackage{enumitem}

\title{LEGAL-UQA: A Low-Resource Urdu-English Dataset for Legal Question Answering}

\author{
  \textbf{Faizan Faisal}\thanks{Equal contribution by authors.} \\
  LUMS \\
  Lahore, Pakistan \\
  \texttt{23100030@lums.edu.pk} \\\And
  \textbf{Umair Yousaf}\footnotemark[1] \\
  LUMS \\
  Lahore, Pakistan \\
  \texttt{23100053@lums.edu.pk} \\
}

\begin{document}
\maketitle

\begin{abstract}

We present LEGAL-UQA, the first Urdu legal question-answering dataset derived from Pakistan's constitution. This parallel English-Urdu dataset includes 619 question-answer pairs, each with corresponding legal article contexts, addressing the need for domain-specific NLP resources in low-resource languages. We describe the dataset creation process, including OCR extraction, manual refinement, and GPT-4-assisted translation and generation of QA pairs. Our experiments evaluate the latest generalist language and embedding models on LEGAL-UQA, with Claude-3.5-Sonnet achieving 99.19\% human-evaluated accuracy. We fine-tune mt5-large-UQA-1.0, highlighting the challenges of adapting multilingual models to specialized domains. Additionally, we assess retrieval performance, finding OpenAI's text-embedding-3-large outperforms Mistral's mistral-embed. LEGAL-UQA bridges the gap between global NLP advancements and localized applications, particularly in constitutional law, and lays the foundation for improved legal information access in Pakistan.
\end{abstract}

\section{Introduction}

While Natural Language Processing (NLP) has seen major advancements in recent years, the benefits of these advancements remain disproportionately skewed towards a small subset of global languages, specifically English \cite{electronics13030648}. This linguistic imbalance results in significant accessibility gaps, particularly in critical domains like law.

Urdu is Pakistan's national language and currently the tenth most spoken language in the world when combining first and second language speakers \cite{cia_world_factbook}. It is the first language of over 2.2 million people in Pakistan \cite{pbs_population_2023}. While English serves as Pakistan's official language alongside Urdu, most government, legal, and administrative documentation is conducted in English. However, only a small fraction of judges, lawyers, litigants, and defendants are proficient in English for court proceedings. This limited English proficiency often hinders justice, particularly for marginalized groups, especially women \cite{ahmad2020english}. Recognizing this, the Pakistani constitution states:

\begin{quote}
    "English language may be used for official purposes until arrangements are made for its replacement by Urdu." \cite{constitution_of_pakistan_english}
\end{quote}

Pakistan's technology landscape is currently lagging, particularly in the legal domain. Studies such as \cite{khan_bridging_2024} highlight how legal chatbots can provide instant, round-the-clock guidance on basic legal queries, procedures, and rights, helping users navigate legal processes and understand legal jargon.

We propose the first Urdu-based legal Question-Answering (QA) dataset, derived from the constitutional law of Pakistan. We aim to develop a resource that is both foundational and broadly applicable, covering essential legal rights and obligations that are of public interest. The dataset is a closed-domain Generative QA dataset, where the abstractive answer to a question is generated by a generative model, in contrast to an Extractive QA dataset, where the answer is directly extracted from the provided context \cite{wang2022modernquestionansweringdatasets}. We open-source all our code\footnote{\url{https://github.com/nlp-anonymous-researcher/LEGAL-UQA/tree/main}}, dataset\footnote{\url{https://huggingface.co/datasets/nlp-anonymous-researcher/LEGAL-UQA}}, and fine-tuned model\footnote{\url{https://huggingface.co/nlp-anonymous-researcher/mt5-large-legal-uqa/tree/main}}.

Our contributions are threefold:
\begin{enumerate}[itemsep=0pt]
    \item \textbf{Dataset Creation}: We introduce a parallel QA dataset containing questions, answers, and their corresponding relevant context (each law article present in the constitution). It is designed to support machine learning models in understanding legal queries in Urdu. Article-wise chunking in the dataset also opens doors for Retrieval-Augmented Generation (RAG)-based legal chatbots \cite{NEURIPS2020_6b493230}.
    
    \item \textbf{Domain-Specific Focus and Linguistic Relevance}: As discussed later in Section~\ref{sec:related_work}, while large Urdu QA datasets exist, domain-specific datasets are scarce. We contribute to the growing body of work in multilingual NLP, offering a resource that bridges the gap between global NLP research and localized, culturally relevant applications.
    
    \item \textbf{Experimentation with LLMs}: Our study evaluates the dataset using a variety of the latest Large Language Models (LLMs). Additionally, we fine-tune mt5-large-UQA-1.0 \cite{arif-etal-2024-uqa} to compare its results with those of generalist models.
\end{enumerate}

\section{Related Work}
\label{sec:related_work}
While there exist several Urdu question-and-answer (QA) datasets, to the best of our knowledge, none are adapted for the legal domain, resulting in a significant gap in this sector. Finding diverse datasets for low-resource languages is challenging, and a common technique involves translating existing datasets into the required language.

\subsection{Urdu QA Datasets}
UQA \cite{arif-etal-2024-uqa}, introduced in May 2024, is built on SQuAD2.0 \cite{rajpurkar_know_2018} and contains $\sim$88k answerable questions and $\sim$47k non-answerable questions. The authors used seamless M4T \cite{communication_seamlessm4t:_2023} to translate the question-answer pairs.

UQuAD (Urdu Question-Answer Dataset)\footnote{\url{https://github.com/ahsanfarooqui/UQuAD---Urdu-Question-Answer-Dataset}} includes 27 paragraphs and 499 question-answer pairs, and is available on GitHub.

UQuAD1.0 \cite{kazi_uquad1.0:_2021} contains $\sim$46.5k questions across $\sim$18.8k paragraphs, derived from a range of Wikipedia articles. These articles are distributed across various categories such as politics, religion, education, music, and miscellaneous. This dataset was created through the machine translation of SQuAD1.0 \cite{rajpurkar-etal-2016-squad}.

The Urdu Open\footnote{\url{https://www.futurebeeai.com/dataset/prompt-response-dataset/urdu-open-ended-question-answer-text-dataset}} and Close-Ended\footnote{\url{https://www.futurebeeai.com/dataset/prompt-response-dataset/urdu-closed-ended-question-answer-text-dataset}} Question Answer Text Datasets are two additional QA Datasets. The open-ended dataset contains questions with corresponding answers in Urdu, without any predefined context. These cover topics such as science, history, technology, literature, and current affairs, and were compiled by native Urdu speakers from reliable sources, including books, articles, and websites. In contrast, the closed-ended dataset includes question-answer pairs accompanied by context. This dataset also draws from reliable sources. However, both datasets are not publicly available, and no sample data is provided.

\subsection{Legal QA Datasets}
In the legal QA domain, several datasets exist. The JEC-QA Legal QA dataset \cite{zhong_jec-qa:_2019}, created from China's National Judicial Examination, contains 26,365 multiple-choice and multiple-answer questions. Its objective is to predict answers based on questions and relevant legal articles, and it is in Chinese.

LeDQA \cite{liu_ledqa_2024}, another Chinese dataset, includes legal case documents with corresponding QA pairs that can be answered using the document as context. It comprises 100 case documents, 4,800 case-question-answer triplets, and 132,048 sentence-level relevance annotations. Models such as Qwen-7b-chat \cite{bai_qwen_2023}, GPT-3.5-turbo\footnote{\url{https://platform.openai.com/docs/models/gpt-3-5-turbo}}, and ChatLaw \cite{cui_chatlaw:_2023} have been tested on this dataset.

EQUALS \cite{EQUALS} contains 6,914 (question, article, answer) triplets, along with a pool of law articles covering 10 Chinese law collections. The questions and answers are sourced from a legal consultation forum, with law article spans annotated by senior law students, ensuring high quality.

\begin{table*}[t]
    \centering
    \caption{Statistics of Article Categories used to create English QA Pairs}
    \label{tab:category_stats}
    \begin{tabular}{lccccc}
        \toprule
        \textbf{Category} & \textbf{No. of Articles} & \textbf{No. of Questions} & \textbf{Mean Word Count} & \textbf{Std. Dev.} & \textbf{Range} \\
        \midrule
        Small & 79 & 79 & 31.05 & 8.21 & [8, 43] \\
        Medium & 154 & 308 & 112.40 & 49.11 & [45, 233] \\
        Large & 79 & 237 & 438.68 & 224.87 & [235, 1336] \\
        \bottomrule
    \end{tabular}
\end{table*}

\section{Dataset Description}

\subsection{Document Parsing for Contexts}
We used the official document of the Constitution, which is available in English on the official government website for the National Assembly \cite{constitution_of_pakistan_english}. While the English version contained text, the Urdu version consisted of images \cite{constitution_of_pakistan_urdu}. We employed the current state-of-the-art Urdu OCR, UTRNet \cite{rahman_utrnet_2023}, available on GitHub\footnote{\url{https://github.com/abdur75648/UTRNet-High-Resolution-Urdu-Text-Recognition}}, to extract the text from the Urdu images. UTRNet proposed bounding boxes for lines, and \cite{rahman_utrnet_2023} provided model weights for YoloV8 \cite{10533619} trained on UrduDoc \cite{rahman_utrnet_2023}. This model was used to carry out text detection on each proposed line to generate the final texts.

On manually inspecting the results of the OCR, we noticed occasional formatting errors. Spelling errors and jumbled sentences were rare. To rectify these errors made by the OCR, we pass each of the parsed articles in Urdu to GPT-4o \cite{openai_models} with the prompt in Appendix~\ref{sec:appendixB}. This helped to minimize any occasional incorrect formatting, spelling errors, and jumbled sentences.

After both documents were converted to text-based PDF format, we transformed them into editable Word documents to manually insert delimiters (triple backticks) at the start and end of each of the 312 legal articles in both documents. We first tested automated techniques for parsing articles, such as Unstructured.io \footnote{\url{https://github.com/Unstructured-IO/unstructured}} and LLM Sherpa \footnote{\url{https://github.com/nlmatics/llmsherpa}}, but they produced errors in correctly classifying headings and body text.


\begin{table*}[h]
    \centering
    \caption{QA Model Evaluation Results}
    \begin{tabular}{lcccc}
        \toprule
        \textbf{Model} & \textbf{F1} & \textbf{METEOR} & \textbf{SacreBLEU} & \textbf{Accuracy (Human)} \\
        \midrule
        Claude-3.5-Sonnet & \textbf{66.57} & \textbf{0.689} & 44.87 & \textbf{99.19} \\ 
        Gemini-1.5-Flash-002 & 35.17 & 0.276 & 18.00 & 98.39 \\ 
        Gemini-1.5-Pro-002 & 41.32 & 0.341 & 24.60 & 98.39 \\ 
        Mistral-Large & 65.91 & 0.614 & \textbf{51.45} & 96.77 \\ 
        mt5-large-UQA-1.0 & 53.19 & 0.437 & 21.23 & 74.19 \\ 
        \bottomrule
    \end{tabular}
    \label{tab:qa_model_eval}
\end{table*}

\begin{table}[h]
    \centering
    \caption{Top-k retrieval results using Cosine Similarity}
    \label{tab:retrieval_results}
    \begin{tabular}{lccc}
        \toprule
        \parbox{2.5cm}{\textbf{Embedding Model}} & \textbf{k=1} & \textbf{k=3} & \textbf{k=5} \\
        \midrule
        \parbox{2.5cm}{text-embedding-3-large} & \textbf{53.23\%} & \textbf{73.39\%} & \textbf{77.42\%} \\
        mistral-embed & 29.03\% & 45.97\% & 52.42\% \\
        \bottomrule
    \end{tabular}
\end{table} 

\subsection{English QA Pairs}

To determine the number of questions to generate from each article, we classified the English articles into three categories—small, medium, and large—based on word count. The bottom 25\% were classified as small, and the top 25\% as large. Table~\ref{tab:category_stats} summarizes the statistics for these categories.

To create a dataset consisting of question-answer (QA) pairs, we utilized OpenAI's GPT-4 model \cite{openai_models}. Short articles were used to generate a single QA pair, whereas medium- and large-sized articles were used to generate two and three QA pairs, respectively. The detailed prompt is available in the Appendix ~\ref{sec:appendixA}. In each iteration, the model received an article and generated QA pairs in English related to its content. We refined the prompt for a more natural tone, generating a total of 619 QA pairs after manually filtering out three invalid questions and two duplicates.

\subsection{Urdu QA Pairs}
To generate Urdu versions of the English QA pairs, we used OpenAI's GPT-4o model, providing it with an English context (a single English article), a corresponding English QA pair and Urdu context. The Urdu context was provided to ground the model in the Urdu language style used in the official Urdu Constitution document. The model was prompted to translate the English QA pair while maintaining the style and tone of the Urdu context. The detailed prompt can be viewed in the Appendix~\ref{sec:appendixC}. We randomly sampled 30 Urdu QA pairs to ensure that the translations are valid.

\subsection{Post-Processing}
We combined the features of the dataset, which consists of 619 QA pairs and their corresponding contexts. The features are as follows:
\begin{itemize}[itemsep=0pt]
    \item \texttt{question\_eng}: The question in English.
    \item \texttt{question\_urdu}: The question in Urdu.
    \item \texttt{context\_eng}: The context in English.
    \item \texttt{context\_urdu}: The context in Urdu.
    \item \texttt{answer\_eng}: The answer in English.
    \item \texttt{answer\_urdu}: The answer in Urdu.
    \item \texttt{context\_index}: A unique identifier for each context/article.
\end{itemize}

This is a Generative QA dataset, as opposed to an Extractive QA dataset, where answer spans are included, and answers are exact phrases from the provided context \cite{wang2022modernquestionansweringdatasets}.

\section{Experiments \& Evaluation}

We conducted experiments on several generative LLMs. We created a validation set consisting of 124 QA pairs and a training set with 495 data points. We fine-tuned the mt5-large-UQA-1.0 model on our dataset. This model is a fine-tuned version of mt5-large \cite{xue-etal-2021-mt5}, trained on the UQA corpus. We used a single NVIDIA A100 GPU on Google Colab\footnote{\url{https://colab.research.google.com/}}. 
We used a learning rate of 5e-5 and fine-tuned the model for 10 epochs, achieving the lowest validation accuracy in epoch 2, whose weights were used in all subsequent experiments.

Furthermore, we conducted experiments with the latest models from Google \cite{google_gemini}, Anthropic \cite{anthropic-docs-models}, and Mistral \cite{mistral_embeddings}. We did not evaluate OpenAI models to avoid any potential bias, as they are used to augment the dataset.

To evaluate the QA results, we used F1 \cite{rajpurkar-etal-2016-squad}, METEOR \cite{banerjee-lavie-2005-meteor}, and SacreBLEU \cite{post2018clarityreportingbleuscores}. Additionally, the accuracy of the answers was manually assessed. The answers were marked as correct (1) or incorrect (0) based on whether the question was accurately answered. 

Claude-3.5-Sonnet \cite{anthropic-docs-models} outperformed the other models with a human accuracy of 99.19\% (Table \ref{tab:qa_model_eval}). Gemini \cite{google_gemini} models exhibited lower performance on the metrics despite achieving high human accuracy. While Gemini's outputs were correct, they were extremely concise compared to those in our dataset, hence the discrepancy. The fine-tuned mt5-large-UQA-1.0 model struggled to answer questions and frequently made common grammatical errors.

To evaluate retrieval performance, we generated embeddings for the validation set questions and the combined training and validation contexts. For each question, the top-k closest embeddings were identified using cosine similarity \cite{han2012data}. A score of 1 was given if the relevant context was found in the closest matches; otherwise, 0.

As shown in Table \ref{tab:retrieval_results}, OpenAI's text-embedding-3-large \cite{openai_models} outperformed all other models in retrievals, while Mistral's mistral-embed \cite{mistral_embeddings} performed poorly. Google's text-embedding-004 \cite{google_gemini} yielded disappointing results due to its focus on English-only embeddings. Access to their text-multilingual-embedding-002 model \cite{google_vertex_ai_embeddings} was not obtained. The open-source Urdu embedding model mt5-base-finetuned-urdu\footnote{\url{https://huggingface.co/eslamxm/mt5-base-finetuned-urdu}} was also evaluated but showed extremely poor performance.

\section{Conclusion and Future Work}
Our study introduces LEGAL-UQA, the first legal QA dataset in Urdu, derived from Pakistan's constitution.

Experiments, particularly with the mT5 model, reveal challenges for low-resource languages in specialized domains. We plan to expand the dataset to include criminal, civil, and administrative law, enhancing its applicability. Additionally, we aim to explore more open-source models. The dataset's parallel nature could also improve legal translations.

These initiatives seek to enhance access to legal information for Urdu speakers in Pakistan and advance multilingual NLP for low-resource languages in the legal domain.

\section{Limitations}
The LEGAL-UQA dataset, while the first of its kind, consists of a relatively small set of 619 QA pairs in the context of LLMs. Although the dataset covers Pakistan's constitutional law, it does not encompass the broader legal landscape of criminal or civil law, limiting its applicability. Additionally, the use of GPT-4 to generate the QA pairs may introduce subtle errors in style or tone.

Another limitation is the reliance on pre-trained models and embeddings that are not optimized for legal text, particularly in low-resource languages like Urdu. Although we fine-tuned the mT5 model, its overall performance lagged behind other state-of-the-art models. Additionally, the evaluators, while fluent in Urdu, had limited legal expertise.

\section{Ethical Considerations}
Building legal datasets can pose risks of misinterpretation by machine learning models. This dataset is intended to assist and promote research in the area and does not aim to replace legal counsel.

\bibliography{references}

\clearpage

\appendix

\section{QA Pair Generation}
\label{sec:appendixA}
The following prompts were used to generate English QA pairs:

\begin{tcolorbox}[title=System Prompt, colback=white, colframe=black, fonttitle=\bfseries]
\raggedright
\small
\ttfamily
You are an expert in interpreting Pakistani legal documents. 
Given an article text from a legal document, you generate a set of questions and their corresponding answers.
\end{tcolorbox}

\begin{tcolorbox}[title=Task Prompt, colback=white, colframe=black, fonttitle=\bfseries]
\raggedright
\small
\ttfamily
Your task is to generate up to \{number\_of\_questions\} unique question(s) and answer(s) based on the following article text: \\[1em]

Article Text:
\texttt{```}\{article\_text\}\texttt{```} \\[1em]

Instructions: \\
- If it is not possible to generate \{number\_of\_questions\} question(s), provide as many questions and answers as you can based on the content. \\
- If no question can be generated, respond with 'NONE'. \\
- Ensure that each question is unique and related to the text provided. \\
- The answer should be based on the information provided in the article text only. \\
- Do not mention article numbers in the questions. \\
- Do not mention the article text in the questions. \\
- You are to ask questions as a layman would ask naturally, not as a legal expert. They should be from the perspective of a person who is trying to find out what the constitution says about a particular topic. \\
- The questions should not be too specific or too general. \\
- Avoid using the same words or phrases from the article text in the questions. \\
- The answers can use the same words or phrases from the article text. \\[1em]

Example Output Format: \\
Question: What is the official name of the country Pakistan? \\
Answer: The official name of the country is the Islamic Republic of Pakistan. \\[1em]

Question: Which regions are part of Pakistan? \\
Answer: The regions mentioned as parts of Pakistan are Balochistan, Khyber Pakhtunkhwa, Punjab, Sindh, and the Islamabad Capital Territory.
\end{tcolorbox}

\noindent
Where:
\begin{enumerate}
    \item \texttt{\{number\_of\_questions\}} is the number of QA pairs to be generated
    \item \texttt{\{article\_text\}} is the context in English
\end{enumerate}

\section{Refining Urdu OCR Detections}
\label{sec:appendixB}
The following prompts were used on the results of the end-to-end Urdu OCR pipeline to refine the detections.

\begin{tcolorbox}[title=System Prompt, colback=white, colframe=black, fonttitle=\bfseries]
\raggedright
\small
\ttfamily
You are an expert in both English and Urdu languages.
\end{tcolorbox}

\begin{tcolorbox}[title=Task Prompt, colback=white, colframe=black, fonttitle=\bfseries]
\raggedright
\small
\ttfamily
You are given an English text and its Urdu translation, which contains formatting and grammatical errors. \\ [1em]
Your task is to:
\begin{enumerate}
    \item Correct any grammatical or formatting mistakes in the Urdu text.
    \item Properly format bullet points or lists.
    \item Re-order sentences if they appear jumbled.
\end{enumerate}

Output Instructions: \\
Only provide the corrected Urdu text without any explanations or comments. \\[1em]

English Text: \\
\texttt{```}\{article\_text\_eng\}\texttt{```} \\[1em]
    
Urdu Text: \\
\{article\_text\_urdu\}
\end{tcolorbox}

\noindent
Where:
\begin{enumerate}
    \item \texttt{\{article\_text\_eng\}} is the original English context
    \item \texttt{\{article\_text\_urdu\}} is the unrefined Urdu context
\end{enumerate}

\clearpage

\section{Translation}
\label{sec:appendixC}

The following prompt instruction was used in our study for translation of English QA pairs to Urdu:

\begin{tcolorbox}[title=Translation Prompt, colback=white, colframe=black, fonttitle=\bfseries]
\raggedright
\small
\ttfamily
You are provided with the **context in English** as well as the **Q\&A pair in English**. You will also be given the **context in Urdu**. \\[1em]

Your task is to translate the English Q\&A pair into Urdu, ensuring that the translation aligns with the style and tone of the Urdu context provided.\\[1em]

**Context in English:** \{context\_en\} \\[1em]

**Question (in English):** \{question\} \\[1em]

**Answer (in English):** \{answer\} \\[1em]

**Context in Urdu:** \{context\_ur\} \\[1em]

\{format\_instructions\}
\end{tcolorbox}

\noindent
Where:
\begin{enumerate}
    \item \texttt{\{context\_en\}} represents the context provided in English.
    \item \texttt{\{question\}} is the question in English.
    \item \texttt{\{answer\}} is the answer in English.
    \item \texttt{\{context\_ur\}} represents the context provided in Urdu.
    \item \texttt{\{format\_instructions\}} are additional pydantic formatting instructions given to the model to parse the output.
\end{enumerate}

\end{document}